There were 90 children at the birthday party.If 3/5 of them were boys,how manay were girls? Step by step
Found 2 solutions by t0hierry, rothauserc:
Answer by t0hierry(194)

3/5 of 90 children were boys. That means 2/5 of 90 were girls, because 3/5 + 2/5 = 5/5 = 1.
2/5 * 90 = 180/5 = 36
Therefore there were 36 girls at the birthday party.

Answer by rothauserc(4718)

90 children at the birthday party
:
90 * (3/5) = 54 of the children are boys, then
:
90 - 54 = 36 of the children are girls
: